OpenStudy (anonymous):

PLEASE HELP :)))) Simplify the expression. 1-cot x/ tan x-1 cot x tan x -cot x csc x

OpenStudy (mertsj):

\[\frac{1-\frac{cosx}{sinx}}{\frac{sinx}{cosx}-1}=\]

OpenStudy (mertsj):

\[\frac{\frac{sinx-cosx}{sinx}}{\frac{sinx-cosx}{cosx}}=\]

OpenStudy (mertsj):

\[\frac{sinx-cosx}{sinx}\times \frac{cosx}{sinx-cosx}=\frac{cosx}{sinx}=cotx\]

OpenStudy (anonymous):

Thank you!!!

OpenStudy (mertsj):

yw